OpenStudy (pphalke):

-7^2 how do I do this?

OpenStudy (vhtran):

you can always put it into a calculator

OpenStudy (vhtran):

its basically saying \[-7 \times -7\]

OpenStudy (anonymous):

if the power is 2, 4 and so on, the - dissapears. And you just square the 7 . Or -7 to the power of 2 is basically (-7 * -7), for example 2 to the power of 3 would be ( 2*2*2)

OpenStudy (skullpatrol):

$$\huge -7^2=-(7•7)$$

OpenStudy (vhtran):

try putting it into a calculator..it should come out as 49, a negative multiplied by another negative gives you a positive. \[-7 \times -7=49\]

OpenStudy (skullpatrol):

Think of $$-x^2$$ with 7 in place of x.

OpenStudy (vhtran):

that would be the case if it was \[-(7^2)\]

OpenStudy (skullpatrol):

$$ y=-x^2$$ is a parabola that opens downward because of the minus sign.

OpenStudy (vhtran):

just...put it into a claculator

OpenStudy (skullpatrol):

A good way to read -7^2 is "the opposite of the square of 7" and to recall the order of operations